If the pressure at sea level is 101.325 kPa, and assuming a constant density of 1.2550 kg/m3 , what will be the pressure at an altitude of 11,000 m


1
Expert's answer
2022-05-05T02:19:02-0400
"p=p_0e^{-\frac{Mgh}{RT}}=p_0e^{-\frac{g\rho_0h}{p_0}}\\
=101325e^{-\frac{9.8*1.2550*11000}{101325}}=26659\:\rm Pa"